OpenStudy (anonymous):

A salesman had to drive from his home to meet a client. If he drove 1/4 of the distance on Monday and 1/3 of the distance on Tuesday, what portion of the trip is left for him to travel on Wednesday?

OpenStudy (anonymous):

change to common denominator and add the two fraction

OpenStudy (lgbasallote):

let x be the total distance 1/4 x is what he drove on monday 1/3 x is what he drove on tuesday so the remaining distance would be \[\implies x - \frac 14 x + \frac 13x\]

OpenStudy (lgbasallote):

wait...that should be \[\implies x - \left (\frac 14 x + \frac 13 x\right )\] sorry bout that

OpenStudy (anonymous):

Thank you so much bro

OpenStudy (anonymous):

I'm still having trouble

OpenStudy (lgbasallote):

where?

OpenStudy (anonymous):

Imm getting 1/6 but it's wrong

OpenStudy (lgbasallote):

first tell me.. what did you get for 1/4 x + 1/3 x

OpenStudy (anonymous):

2/12?

OpenStudy (lgbasallote):

\[\frac 14 + \frac 13 \implies \frac{3}{12} + \frac{4}{12} = \frac{7}{12}\] got it?

OpenStudy (anonymous):

I think but how did u get the 3/12 + 4/12?

OpenStudy (lgbasallote):

it's changing to LCD

OpenStudy (lgbasallote):

1/4 you'll change the denominator to 12 so 12 divided by 4 is 3 then 3 times 1 is 3 so 3/12

OpenStudy (lgbasallote):

it's hard to explain in words =))

OpenStudy (anonymous):

I think I got it though..thanks a lot u were a big help!!

OpenStudy (lgbasallote):

welcome ^_^